# The Race to Efficiency: A New Perspective on AI Scaling Laws

### v2

@@ 3.1 Key Parameters and Notation
    Logical (Model) FLOPs, defined to remain stable, vendor-agnostic, and consistent with industry standards;
@@ 5.2 A Multi-Year Case: Baseline, Turtle, and Hare
    Smaller (Possibly Older) GPU Fleet: Starting with fewer GPUs (e.g., 2,000 rather than 10,000) and aiming for a higher efficiency-doubling rate ($\gamma>2$) to offset the lower initial capacity.
@@ 5.4 Outlook
Recent NVIDIA data shows a 45,000$\times$ leap in token-generation efficiency over eight years—effectively doubling every six months ($\gamma\approx 2$[20]. Meanwhile, organizations like OpenAI, Google, and Meta continuously refine hardware, software, data pipelines, and infrastructure to maintain similarly rapid efficiency-doubling rates. For instance, DeepSeek-V3 [18] achieves 17$\times$ higher efficiency (see Appendix A) compared to Llama 3 (405B) through innovations such as FP8 arithmetic, sparsity, and Mixture-of-Experts, demonstrating how architectural optimizations can drive significant efficiency gains. The success of DeepSeek-V3 underscores the importance of prioritizing efficiency improvements in AI development, particularly for organizations with limited compute budgets or sustainability goals. This $\gamma>0$ trajectory is far from theoretical: it underpins the near-exponential performance gains needed for training modern large-scale models.

### v3

@@ Abstract
[…]
The Race to Efficiency: A New Perspective on AI Scaling Laws
Chien-Ping Lu
cplu@nimbyss.com
@@ 3.1 Key Parameters and Notation
    Logical (Model) FLOPs, defined to remain stable, vendor-agnostic, and consistent with industry standards of comparing ”teraflops” among different precisions such as FP16, BF16, FP8, and FP4;
@@ 5.2 A Multi-Year Case: Baseline, Turtle, and Hare
    Smaller GPU Fleet: Starting with fewer GPUs (e.g., 2,000 rather than 10,000) and aiming for a higher efficiency-doubling rate ($\gamma>2$) to offset the lower initial capacity.
@@ 5.4 Outlook
Recent industry data highlights significant advancements in AI energy efficiency, with NVIDIA reporting a 45,000$\times$ improvement in energy efficiency for AI inference over the past eight years—equating to doubling approximately every six months ($\gamma\approx 2$[20]. This progress is exemplified by the latest GB10, which integrates the Grace CPU and Blackwell GPU into a desktop form factor [21]. Concurrently, organizations like OpenAI, Google, and Meta continue to refine hardware, software, data pipelines, and infrastructure to sustain rapid improvements in efficiency. For example, DeepSeek-V3 [18] achieves substantial efficiency gains through innovations such as FP8 arithmetic, sparsity, and Mixture-of-Experts, demonstrating the potential of architectural optimizations. The success of platforms like Grace Blackwell and innovations such as DeepSeek-V3 underscore the critical importance of prioritizing efficiency improvements in AI development, particularly for organizations with constrained compute resources or sustainability targets. This sustained trajectory of efficiency improvement directly supports the exponential performance gains necessary for training modern large-scale models.